OpenStudy (anonymous):

lim x->0 sin4x/x , can somone help me solve this step by step ?? thanks

OpenStudy (anonymous):

sorry i mean sin4x/7x

OpenStudy (zarkon):

\[\lim_{x\to 0}\frac{\sin(4x)}{7x}=\lim_{x\to 0}\frac{4}{7}\frac{\sin(4x)}{4x}\]

OpenStudy (anonymous):

ok then ??

OpenStudy (zarkon):

do you know this limit \[\lim_{x\to 0}\frac{\sin(x)}{x}\]

OpenStudy (anonymous):

yes that equals to 1 , so i am assuming the answer is 4/7?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

how should i write it ?

OpenStudy (anonymous):

could u contunie the steps , becuz my teacher is a hard marker

OpenStudy (zarkon):

\[\lim_{x\to 0}\frac{4}{7}\frac{\sin(4x)}{4x}\] let \(t=4x\) then as \(x\to 0\) we have \(t\to 0\) and thus \[\lim_{x\to 0}\frac{4}{7}\frac{\sin(4x)}{4x}=\lim_{t\to 0}\frac{4}{7}\frac{\sin(t)}{t}=\frac{4}{7}\times 1\] \[=\frac{4}{7}\]

OpenStudy (anonymous):

could u help me solve one more??

OpenStudy (anonymous):

tanx/3x

OpenStudy (zarkon):

same idea \[\frac{\tan(x)}{3x}=\frac{\sin(x)/\cos(x)}{3x}=\frac{\sin(x)}{3x}\frac{1}{\cos(x)}\] now do what I did above

OpenStudy (zarkon):

\[=\frac{1}{3}\frac{\sin(x)}{x}\frac{1}{\cos(x)}\]

OpenStudy (anonymous):

i know the fist funstion equals 1 what about 1/cosx?

OpenStudy (anonymous):

@Zarkon

OpenStudy (zarkon):

is \(x\to 0\)?

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

what is \(\cos(0)\)?

OpenStudy (anonymous):

1?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

so their both equal to 1 ?? and so the answer is 1/3??

OpenStudy (zarkon):

yes